OpenStudy (anonymous):

The equation of a line is x + 5y = 12. What is the y-intercept of the line? Option A: -12 Option B: 12 Option C: 5/12 Option D: 12/5

OpenStudy (stefrheart):

first you need to move isolate the y intercept onto one side

OpenStudy (anonymous):

would it be 12?

OpenStudy (texaschic101):

put your equation in y = mx + b form, and the b position is the y intercept

OpenStudy (anonymous):

go on mathway.com it could help :)

OpenStudy (anonymous):

I think it's 12 but I'm not sure

OpenStudy (texaschic101):

no..its not 12

OpenStudy (texaschic101):

put your equation in y = mx + b form x + 5y = 12 --- subtract x from both sides 5y = -x + 12 -- divide both sides by 5 y -x/5 + 12/5 now..in y = mx + b form, the y intercept is in the b position. So what do you think the answer is now ?

OpenStudy (texaschic101):

oops...typo...y = -x/5 + 12/5

OpenStudy (anonymous):

12/5?

OpenStudy (anonymous):

D

OpenStudy (texaschic101):

you got it...12/5 is the y intercept. And just to let you know, if you were looking for the slope, it is in the m position...it would be -1/5

OpenStudy (anonymous):

oh yeaah, I get it now. Thank you sooooo much! :))

OpenStudy (texaschic101):

anytime :)